# Applications and resource estimates for open system simulation on a quantum computer

### v1

@@ V.2 Compiled Logical
The weight of Bravyi-Kitaev [bravyi2002fermionic] creation operator goes logarithmic with the system size $10\times 10$ in the MIT example, but we will treat it as a small constant. In other words, we will assume that each $L_{\mu}$ is local in what follows.
@@
Recall that the algorithm in Theorem III.2 of [chen2023quantum] uses the following number of consecutive calls to the block encoding $U$:
@@
where $\alpha$ is the time upscaling factor discussed above ($\alpha=3000$ and $2000$ for the two cases we consider) . To estimate the $T$-gate count, we multiply it by the cost of the block encoding (we neglect the increase in $T$-count due to the controlled use of $U$), and drop the constabt factor hidden in the big-O notation:
#$\#T=\alpha t\frac{\text{log}(\alpha t/\epsilon)}{\text{loglog}(\alpha t/\epsilon)}10M\text{log}M\ ,$ (17)
and same with loglog $M$ for the depth. For the MIT material application $t=54000$ and $\epsilon=0.1$, while for Ca3Co2O6, we need to makes $10$ steps of $t=10^{11}$, each with precision $\epsilon=0.01$. Plugging in these numbers, we get:
MIT search Ca3Co2O6
$T$-count $2.7\cdot 10^{14}$ $1.9\cdot 10^{23}$
$T$-depth $7.0\cdot 10^{13}$ $4.0\cdot 10^{22}$
We see that the parallelism in $T$-gate application is not substantial in this estimate, accounting for only ×$\times 1/4$ reduction in runtime compared to applying $T$-gates one at a time. In other words, the algorithm consumes $4$ $T$ gates per gate layer on average. Ideally, ×1$\times$ 1/× 1 /number of qubits reduction should be possible, yet the QSP algorithms we use do not allow it. We also note that for the MIT material application, the Hamiltonian part of the evolution (not estimated here) may, in fact, be the leading contribution.
V.3 Physical
To estimate the physical requirements, we follow the resource estimate in [beverland2022assessing, Table IV]. The $T$-count can be found in the column $M$, and $(ns,10^{-3})$ is the closest relevant noise parameter for superconducting technology in that reference. We first see that the runtime is roughly $10^{-6}$s times the $T$-count across the application instances they consider.
We multiply the $T$-gate counts by $10^{-6}$s to estimate the total runtime. Since Ca3Co2O6 only needs to be executed once, that results in $1.9\cdot 10^{17}$ seconds ($6\cdot 10^{9}$ yrs), an unrealistic number. Each run of MIT material search requires only $2.7\cdot 10^{8}$s (8.6 years). We will be able to reduce that number at the cost of increasing the number of qubits below.
A more careful resource estimate for the physical count would proceed as follows. We will follow a recent short paper [hirano2024leveraging]. Specifically, we note that Fig. 17 in it allows one to draw a line through the qubitcycle resource estimate between $10^{-4}$ and $10^{-10}$ error rate of the $T$-state: we get $100,000$ qubitcycles per each power of $10$ below $10^{-4}$ individual $T$-state error. Table II of [hirano2024leveraging] allows us to estimate the time and space requirements: the rightmost cluster of star symbols have $78$ cycles and $10614$ qubits, which gives the ratio of the two in the qubitcycle number. The required error of the distilled state is roughly the target error $\epsilon$/number of $T$ gates. The time of the code cycle is 4 CNOTs and a measurement. In [beverland2022assessing], we find the cycle time from Table II to be $300$ns. The resulting costs per factory become:
Factory time cost $=13(|\text{log}_{10}(\epsilon/\#T)|-4)\text{cycles}=0.4(|\text{log}_{10}(\epsilon/\#T)|-4)\mu$s
Factory qubit cost$=1,800(|\text{log}_{10}(\epsilon/\#T)|-4)$
We will extrapolate beyond $10^{10}$ $T$-gates, though the factory costs may not be the leading cost in that regime. For the two cases we consider, the numbers are:
MIT: 150 cycles, 21,000 qubits, $d\sim 40$
Ca3Co2O6: 280 cycles, 38,000 qubits, $d\sim 50$
The distances are just rough guesses and can be further refined using the surface code threshold formula if necessary. For the entire algorithm, a tradeoff between using multiple factories or using fewer qubits but taking longer is possible. With a single factory, we estimate the total number of physical qubits as the factory size plus $d^{2}$ times the logical number of qubits, and the runtime is
@@ V.3 Physical
The factor ends up $4.4\cdot 10^{-5}$s and $8.3\cdot 10^{-5}$s times the number of $T$-gates for our two applications.
With multiple factories, one of the limiting factors is the number of $T$-gates per layer. If the algorithm consumes on average $k$ $T$-gates per layer (in our case $k=4$), then generating more $T$-states than $k\times$Factory time cost in cycles is not useful as they will just idle until their time in the algorithm comes. So the number of factories $n_{f}$ can vary between 1 and $k\times$Factory time cost in cycles. An additional consideration for this tradeoff is the routing time. We will assume the naive picture where there are pathways between factories, and the throughput of 1 $T$-state per cycle is possible for each pathway. A more careful consideration may restrict this further, thus restricting $n_{f}$. In the case of our two examples, we get:
MIT search Ca3Co2O6
total runtime (cycles) $4\cdot 10^{16}/n_{f}$ $5\cdot 10^{25}/n_{f}$
total runtime (s) $1.2\cdot 10^{10}/n_{f}$ $1.5\cdot 10^{19}/n_{f}$
#factory qubits $21000n_{f}$ $38000n_{f}$
range of $n_{f}$ $1\dots 600$ $1\dots 1100$
runtime (s) for max $n_{f}$ $2.0\cdot 10^{7}$ $1.4\cdot 10^{16}$
($235$ days) ($4\cdot 10^{8}$ yrs)
We see that for $n_{f}=44$ for MIT and $n_{f}=83$ for the Ca3Co2O6, the rule of thumb $10^{-6}$s per $T$-gate is satisfied. We also see that this time cost can be further reduced at the expense of a larger quantum computer. In particular, $6\cdot 10^{9}$ yrs are reduced to $3400$ years on a $42$ million qubit device for Ca3Co2O6 (here, the non-factory $\sim 5$M qubit cost is subleading). For MIT search, each run can now complete in $235$ days using $12$M factory qubits and $\sim 24$M non-factory qubits. We need to execute $10^{3}$ runs for each model, which, if executed in parallel on multiple quantum computers, will give experts time to come up with the next model. We expect only the hardest models, say $1\%$ of 1000 suggested ones, to be sent to the quantum computer, while others investigated by classical methods or experimentally.

### v2

@@ V.2 Compiled Logical
V.2.1 Single Lindblad generators
For Ca3Co2O6, the Lindblad terms (Eq. (6)) contain $\sigma^{\pm}$ multiplied by projectors onto the state of a constant number of neighboring spins. For the Hubbard problem, after the Bravyi-Kitaev [bravyi2002fermionic] or Jordan-Wigner transformation from fermions to qubits, the creation-annihilation operators would have a logarithmic weight, containing a product of Pauli’s and one $\sigma^{\pm}$. We will now briefly outline a possible (not necessarily optimal) block-encoding of those operators. First, we note that $\sigma^{\pm}$ can be block encoded into a 2-qubit (Clifford) permutation:
$U_{+}=X_{0}\,\text{SWAP}=\left(\begin{array}[]{cc|cc}0&1&0&0\\ 0&0&0&1\\ \hline\cr 1&0&0&0\\ 0&0&1&0\end{array}\right)=\left(\begin{array}[]{c|c}\sigma^{+}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right),$ (27)
$U_{-}=X_{1}\,\text{SWAP}=\left(\begin{array}[]{cc|cc}0&0&1&0\\ 1&0&0&0\\ \hline\cr 0&0&0&1\\ 0&1&0&0\end{array}\right)=\left(\begin{array}[]{c|c}\sigma^{-}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right).$ (34)
[…]
The evolution algorithms [chen_quantum_2023] also also require a controlled version $CU_{\pm}$. As counting required T-gates and depth is fairly technical and can distract the reader from our main message, we discuss all the relevant methods for that in App. C, and only quote the results here. $CU_{\pm}$ in particular require 4 T-gates and depth 14.
The block encoding of creation-annihilation operators contains the Pauli’s on system qubits multiplying $U_{\pm}$, which results in the same non-Clifford cost 00 and $4$ for $U_{\mu}$ and $CU_{\mu}$ respectively. The depth is $4$ for $U_{\mu}$ as all the Pauli’s on $\sim 100$ (on average) other qubits can be applied in parallel. For $CU_{\mu}$, one first needs to prepare up to $199$ long register with the repeated value of the control, do every $CZ$ in parallel, and then undo the register. The average depth of that is $\text{log}199\sim 8$ and can done in parallel with the CSWAP. We will later need to control $CU_{\pm}$ on the type information $\pm$. In App. C we show that $C^{2}U_{\pm}$ can be implemented with $12$ T-gates and average depth $37$.
For lower temperatures, we expect to numerically find low-cost block-encodings of Lindblad generators among a family where a local raising or lowering operator is multiplied by functions of the neighboring occupation numbers and terms like $c^{\dagger}_{i}c_{j}$. These terms will only require diagonalization of the system in a local neighborhood. The more complex unit cell of the average case MIT material search requires even more optimization, as there is no notion of locality between orbitals. The straightforward application of the thermal Lindblad generators suggested in [chen_quantum_2023] results in unrealistic (while still polynomial) resource counts, as we show in App. D.
In the Lindblad generator for Ca3Co2O6, instead of Pauli’s, projectors $P_{r\pm}$ on various states of the neighbors multiply $\sigma^{\pm}$. They also need to be block encoded using additional ancillae. Note that since $P_{r\pm}$ is a projection on $14$ spins, one can use one additional ancilla to form a unitary:
$P_{r\pm}\otimes{\mathds{1}}+({\mathds{1}}-P_{r\pm})\otimes X$ (35)
A rank-1 projector, for example, is block-encoded in the following matrix:
$\left(\begin{array}[]{ccccc|ccccc}1&0&\dots&0&0&0&0&\dots&0&0\\ 0&0&\dots&0&0&0&1&\dots&0&0\\ 0&0&\dots&0&0&0&0&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&\dots&0&0&0&0&\dots&0&1\\ \hline\cr 0&0&\dots&0&0&1&0&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&\dots&1&0&0&0&\dots&0&0\\ 0&0&\dots&0&1&0&0&\dots&0&0\\ \end{array}\right)$ (36)
There are multiple ways to implement this unitary, which is essentially a $\tilde{C}X$ conditioned on the states of the qubits to be in the support of $P_{r\pm}$. For a projector rank $k$ with no additional structure, this unitary can be implemented using $k$ applications of $C^{14}X$ (or $2^{14}-k$ applications, whichever is smaller). $k$ varies between $L_{i}$, so we take an average value $2^{14}/39\approx 420$. In other words, it would require $420$ applications of $C^{14}X$. The controlled application would utilize $C^{15}X$ instead. As these numbers are very large, we should instead utilize the structure of $P_{r\pm}$: the spin configurations correspond to particular values of the sum $|J_{1}|\sum_{i\in\text{red}}s_{i}+0.1|J_{1}|\sum_{i\in\text{green,~{}blue}}s_{i}$ of $14$ binary variables, $2$ of which are multiplied by 10. As we didn’t count for possible cancellations, we should simulate the binary adder circuit on groups of $12$ and $2$ and condition the application of $X$ on the result of those additions. The adder is a circuit that prepares a sum of 2 numbers stored as $m$ qubits each in binary. Following [gidney2018halving], when this circuit is used for an operation conditioned on the bits of the sum and then undone, the total cost is $\sim 4m$ T-gates, and the total depth is $\sim 18m$. In our case, instead of adding two numbers, we add single-bit numbers $n=12$ and $n=2$ times, so the largest $m$ needed is $m=3$ for the sum of parts of $12$ spins. No matter how we group them, the total $\sum m=20$, which results in the counts:
$\text{T count}\approx 80,\text{depth}\approx 360$ (37)
As the results of the addition can be stored in 6 qubits, The block encoding of $P_{r\pm}$ also includes $C^{6}X$, and the block encoding of $P_{r\pm}\sigma^{\pm}$ includes additional Cliffords of $\sigma^{\pm}$ that are done in parallel. For $CU$, we will duplicate the ancilla with a CNOT (depth $+2$ for it and its undoing) and apply $C^{7}X$ and CSWAP from $\sigma^{-}$ in parallel. We can translate these gates into T-counts and depths ($C^{7}X\rightarrow 24$ T-gates, depth 57, $C^{6}X\rightarrow 20$ T-gates, depth 53) following App. C. The total counts are:
  • $U_{\mu}$: T-count $100$, depth $413$
  • $CU_{\mu}$: T count $108$, depth $419$
[…]
  • T count: $80$+[C7+kX+ C2+kX count] $=108+8k$
  • depth: $362$+[C7+kX depth]$=393+32\text{log}_{2}(k+1)$
We collect the results as follows:
T-count; depth Hubbard Ca3Co2O6
$U_{\mu}$ 00; $4$ $100$; $418$
$CU_{\mu}$ $4$; $14$ $108$; $427$
$C^{\text{type}}CU_{\mu}$ $12$; $37$ $108+8k$; $393+32\text{log}_{2}(k+1)$
For this table, the type is encoded in the $k=1$ qubit register for Hubbard.
Note that only an $O(1)$ number of ancillae is required for the implementation of the block encoding of the individual Lindblad operators, so it will be neglected in our estimate.
V.2.2 Block-encoding of full Lindbladian
@@
$\text{depth }\sim 16M\text{log}_{2}\text{log}_{2}M$ (39)
Now consider $CU$ that is applied in the time-evolution algorithm. If the system size $n$ is comparable to $M$, it is optimal to add control to every gate in $U$: CNOT-count maps to CCZ-count and T to CT. If in $CU$ several gates controlled on the first qubit need to be applied in parallel, we use a tree of CNOTs to obtain enough classical repetitions of it at the start of $CU$, use them to control the gates in parallel, and undo the CNOTs. This step will only add an $M$-independent constant to depth, which is subleading and will be dropped. The counts for $CU$ are as follows:
[…]
$\text{depth }\sim cM\text{log}_{2}\text{log}_{2}M$ (41)
Here $16\leq c\leq 208$ depending on the positions of CNOTs and Ts.
[…]
We will note that in the case of our applications $L_{\mu}$ is more complex than a single-qubit Pauli for Ca3Co2O6, and for Hubbard $M=20$ is just not large enough, so neither of those problems benefit from the depth reduction. Both, however, strongly benefit from the reduction in the number of circuits for $U_{\mu}$ in the final algorithm that is enabled by this translation invariance trick.
V.3 Translation-invariant block-encoding
First, consider the ancilla $|\mu\rangle$ storing the position information (for now, there is only one type, so the entire $|\mu\rangle$ is used for position). For $M$ different positions, $|\mu\rangle$ encodes the position in binary using $\text{log}M$ qubits. For the first step, we need to map $|\mu\rangle$ into a state on $O(M)$ qubits as follows. For the qubit storing $k$’th bit of position information, we create $M_{k}$ classical repetitions of that bit via a tree of CNOTs. Such a tree has $\sim M_{k}$ CNOTs and depth $\text{log}M_{k}$. The first bit, storing the information about which half of the system the operator is applied to, is repeated $M/2$ times, the next  – $M/4$, and so on, for a total CNOT count $\sim M$ and depth $\sim\text{log}M$ used to generate the classical repetitions for each qubit in $|\mu\rangle$ (on a qubit in a |+ket|$\rangle$| + ⟩ state, this circuit is familiar as the cat state preparation).
We then initiate the SWAPs with the goal to swap the qubit in position $\mu$ to the fixed position $\mu=1$, apply a gate $U_{\mu}$, now not controlled by anything, and swap back. We see that for single-qubit $L_{\mu}$ that are all the same, implementing $\sum_{\mu=1}^{M}|\mu\rangle\langle\mu|\otimes U_{\mu}$ would only require $\sim M$ SWAPs conditioned on one qubit each, and a single deterministic application of $U_{0}$. This is illustrated for $U_{a}=A_{a}$ in Fig. 5.
The leftmost figure of Fig. 5 is the usual way to implement block encoding, which is the typical choice when the unitaries have no structure. In the case of translation invariance, we can, however, rewrite the circuit using only CSWAPs and a single use of $A$ (not controlled). This is essentially an application of binary search; if the ancilla state is |$\ket{\mu}$, to apply $A$ on qubit $\mu$, one can use $O(\log M)$ swaps to move $A$ onto the correct wire. Ancilla bit $k$ has $O(M/2^{k})$ CSWAPs acting. Taking into account all required control swaps means this circuit has depth $O(M)$ and uses $O(\log M)$ ancillae.
This circuit depth can be reduced to $O(\log M)$, and the expense of more ancillae ($O(M)$), which is the rightmost subfigure in 5. The observation is that the swaps in the system register, conditioned on a particular ancilla, act on disjoint pairs of qubits. It is, therefore, possible to first encode the ancilla state using more qubits, allowing application of these CSWAPs in parallel: the $k$’th bit is encoded using CNOTs as |$\ket{\mu_{k}}\rightarrow\ket{\mu_{k}}^{\otimes M_{k}}$, where $M_{k}=M/2^{k}$ (for simplicity, assume $M$ is a power of 2). In Fig. 5 only one CNOT is required, but in general it can be prepared with a circuit of depth $O(\log M_{k})$ [Cruz_2019]. Since each of these GHZ preparation circuits (for each bit of $\mu$) can be implemented in parallel, the depth of this step is $O(\log M)$.
Figure 5: Three equivalent ways of preparing $\sum_{\mu=1}^{M}\ket{\mu}\otimes A_{\mu}\ket{\psi}$, where $A_{\mu}$ is the single qubit unitary $A$ applied on qubit $\mu$. In this example $M=4$. The left/middle figures are implemented with depth $O(M)$ and ancilla cost $O(\log M)$, whereas the rightmost has an ancilla cost $O(M)$ and a depth $O(\log M)$. The curly braces in this figure indicate the CSWAPs can be implemented in parallel. The CNOT comes from the GHZ preparation circuit (see main text).
Thanks to our repeated position register, each of the CSWAPs corresponding to that bit of $|\mu\rangle$ can be done in parallel, so the depth in terms of CSWAPs is $\sim 2\text{log}M$. For the count of $U$, we get:
[…]
$\text{depth of CNOT part}\sim 2\text{log}_{2}M$ (43)
$\text{depth of CSWAP part}\sim 2\text{log}_{2}M$ (44)
In the circuit for $CU$, we again apply control to every gate, assuming the system size is comparable to $M$. The two gates we used are turned into Toffoli and CCSWAP. The resulting counts for $CU$ are:
[…]
ancilla $\text{ancilla count}\geq M$ (46)
We see that while the T-count is comparable with Eq. (41), the T-depth is logarithmic instead of linear, which allows for significant parallelism. The ancilla count depends on the implementation: one can, in principle, undo the repetitions of the first bit and use the same ancilla qubits for the second. In that case, counting also the ancilla in the implementation of CCSWAP, we get ancilla count $\sim M$, thus doubling the qubit count of the algorithm. We also note that if we generalize this approach to a more complicated single-qubit $U_{\mu}$ compiled to multiple T gates, the $M$-independent T-count and depth of $U_{\mu}$ (or $CU_{\mu}$) applied on the qubit at $\mu=1$ is just added to the above counts, thus for sufficiently large $M$ its contribution is subleading.
We will now go over the implications of parallelism during the physical compilation. The fault-tolerant constructions used in the resource estimation [beverland2022assessing] do not allow for T-gate parallelism but benefit from offloading the Clifford depth to classical computation, which also allows to ignore the connectivity constraints. An alternative construction is to keep the Clifford operations and use teleportation for all 2-qubit operations. This will allow for T-gate parallelism but multiply the depth by an additional factor of $5$ for the teleportation depth. The number of logical cycles for $CU$ is then:
[…]
We will see that Ca3Co2O6 still benefits from the translation invariance trick due to the reduction of the number of applications of $U_{\mu}$, but the physical compilation does not need the T-parallelism. For the MIT instance, the special structure of the generators will allow for large parallelism, but our instance is dominated by the Hamiltonian simulation cost.
V.3.1 Cost for Ca3Co2O6
In our applications, unlike the above estimate, $L_{\mu}$ is not a single qubit operator, and there are multiple types of $L_{\mu}$ on the same qubit. For Ca3Co2O6, each spin has $14$ neighbors that need to be swapped together for this trick to work. This severely restricts the depth savings of the trick, as we will illustrate now. This 15-spin set (see Fig. 2) contains 3 red vertices, 6 green and 6 blues, so we do not expect a periodic tiling of the lattice with sets like these (the density of red vertices will be off). There is, however, a periodic tiling where 3 more red vertices are included, say 2 neighbors of the middle one and one neighbor of the bottom one. By slightly translating this tiling 18 possible ways, we get each spin to be the center. This way we can block-encode $\tilde{M}\sim 2025/18\approx 113$ possible translations of each $L_{\mu}$ to other tiles via just one application of the block-encoding of $L_{\mu}$ and the translation trick. $\tilde{M}\sim 113$ is not large enough to justify compiling T gates to execute in parallel, which is why we will instead trade depth for a full translation of all spins. In this approach, the tiling will not be needed. The index $\mu$ will encode one of $2025$ possible shifts of the lattice (we will assume periodic boundary conditions for simplicity). We will now discuss the structure of the circuit $U_{T}$ that shifts a $9\times 9\times 25$ lattice by an arbitrary lattice vector. First note that a sequence of SWAPs $S_{1,2}S_{2,3}S_{3,4}\dots$ shifts the state of the qubits by 1 lattice position. $U_{T}$ can be in principle be implemented by using $2\cdot 2025$ such sequences of length $9$, and $2025$ of length $25$, controlled on an unary encoding of $\mu$. That would cost $\sim 700,000$ T-gates, which is too large. Instead note that a sequence of SWAPs $S_{1,1+k}S_{2,2+k}S_{3,3+k}\dots$ shifts the state by $k$ positions. Any shift of $2^{M}$ qubits can be encoded in an application of $M$ such sequences, with $k=1,2,4\dots 2^{M-1}$. Provided a register $\mu$ encoding the shift, the sequences can be controlled on bits of $\mu$. For our 3d lattice, we need roughly (22025/9)(2$\cdot$ 2025/9$\cdot$( 2 ⋅ 2025 / 9 ) ⋅log9999, (22025/9)(2$\cdot$ 2025/9$\cdot$( 2 ⋅ 2025 / 9 ) ⋅log9999 and (22025/25)(2$\cdot$ 2025/25$\cdot$( 2 ⋅ 2025 / 25 ) ⋅log25252525 sequences of length $9$,$9$ and $25$ respectively. It would cost $178,000$ T gates, and $5000$ depth. There are $39$ types of a unitary $L_{\mu}$ on each spin, each with costs from Sec. V.2.1. The counts of $39$ $C^{k}CU$ with $k=$log39539539$\approx$ 539 ≈ 5 are:
  • T count$=39(108+8k)\sim 7,300$
  • depth $=39(393+32\text{log}_{2}(k+1))\sim 18,500$
The total counts are:
[…]
The leading contributions to the number of qubits are the system and the ancillae controlling SWAPs.
V.3.2 Cost for Hubbard
[…]
There are 40 operators total, so the counts for $U$ are:
[…]
Some savings may be obtained if we note that out of the 40 strings of Jordan-Wigner, 20 has length $<20$, and another 20 has length 180..200. Note that 180 CZ operators are applied conditioned on one bit of the $\mu$ register, thus only require C2Z. The totals are 180 C2Z, $40\cdot 20/2$ C6Z, and $40$ C7Z. The updated counts are:
[…]
We will utilize the same structure for the translation trick: controlled on a single bit of $\mu$, there are two sides with a single application of $\sigma^{\pm}$ respectively and $Z$’s that have to be conditionally applied on up to 19 qubits. The interplay of $Z$ and SWAPs is nontrivial and is illustrated in Fig. 6. A single such circuit on 20 qubits will have $\sim$40 CSWAPs, 19 CZ, and a SWAP for $\sigma^{\pm}$. We need to control it on the additional two qubits: one for the left-right side and another for the control of CU. There are two of those circuits, and 180 C2Z. The counts for CU are:
[…]
The leading contribution to the number of qubits is $3\cdot 200$ for the system and the ancillae controlling SWAPs and $Z$.
V.4 Cost of time evolution
Figure 6: Three equivalent ways of preparing $\sum_{\mu=1}^{M}\ket{\mu}\otimes c_{\mu}^{\dagger}\ket{\psi}$, where $c_{\mu}^{\dagger}$ are creation operators on site $\mu$. In this example $M=4$. The left/middle figures are implemented with depth $O(M)$ and ancilla cost $O(\log M)$, whereas the rightmost has an ancilla cost $O(M)$ and a depth $O(\log M)$. The curly braces in this figure indicate the CSWAPs and CZs can be implemented in parallel. The CNOT comes from the GHZ preparation circuit (see main text).
The T-counts, T-depth, and depth of $CU$ for the two applications can be used to compute parallelism factor $k=T$-count/5/5$\cdot$/ 5 ⋅depth (the factor 5 is due to the depth of teleportation needed for implementing possibly spatially non-local CNOTs in a local architecture), as summarized in a table:
$CU$ Hubbard Ca3Co2O6
T-count $1,200$ $185,000$
qubits $600$ $4000$
depth $32$ $23,000$
$k$ 7.5 1.6
While for Ca3Co2O6 the Hamiltonian is much simpler than the open system generator, and the cost of its block encoding $V$ can be neglected, the opposite is the case for the Hubbard model: while the open system operators are concentrated at the boundaries, the Hamiltonian terms cover the entire lattice. To obtain the counts of the Hubbard Hamiltonian block encoding $V$, we utilize the recently created PyLIQTR library. A brief summary of our numerical experiments is in App. E. The resulting counts are:
[…]
This removes any possible gains of parallelism. It may be possible to recompile the Hamiltonian block encoding using the translation trick and recover the parallelism, which we leave for future work. As there are only 200 qubits, $CV$ can be implemented via CSWAPs of all of them at the cost of $1600$ T gates. The total count for $CV$ and $CU$ is:
[…]
Besides $U$ and $V$, the algorithm in Theorem III.2 of [chen_quantum_2023] also uses other gates. We will assume that the additional 2-qubit gate count $\sim t\text{log}M\text{log}t/\epsilon$ stated in [chen_quantum_2023] can be roughly interpreted as the T-gate count, as it is subleading anyways.
@@ V.4 Cost of time evolution
Recall that the algorithm in Theorem III.2 of [chen_quantum_2023] uses the following number of consecutive calls to $CU$ and $CV$:
@@
where $\alpha$ is the time upscaling factor discussed above ($\alpha=1100$ and $6500$ for the two cases we consider), and $t$ the desired simulation time. To estimate the various counts of the whole algorithm, we multiply it by the corresponding count of the block encoding $CU$ and drop the constant factor hidden in the big-O notation:
@@
For Hubbard, $t=36$ and $\epsilon=0.1$, while for Ca3Co2O6, we need to makes $n_{s}=10$ steps of $t=10^{11}$, each with precision $\epsilon=0.01$. Plugging in these numbers, we get:
Hubbard Ca3Co2O6
$n_{s}\alpha t\frac{\text{log}(\alpha t/\epsilon)}{\text{loglog}(\alpha t/\epsilon)}$ $1.7\cdot 10^{5}$ $6.3\cdot 10^{16}$
T-count $3.3\cdot 10^{9}$ $1.4\cdot 10^{22}$
k 1 1.6
qubits $600$ $4000$
Note that the Trotter open system time evolution algorithm analyzed exactly in App F yields $1.27\cdot 10^{42}$ T-gates for Ca3Co2O6.
V.5 Physical
To estimate the physical requirements, we utilize the Microsoft Azure Resource Estimator introduced in [beverland2022assessing]. We note that recent papers such as [gidney2024magic, hirano2024leveraging] may significantly improve those estimates, especially under the assumption of $10^{-4}$ noise. For our estimates, we use the default $10^{-3}$ noise on superconducting qubits and the AccountForEstimates function with the above numbers. For Ca3Co2O6, $2.1\cdot 10^{22}$ T-gates is too large, and the algorithm fails, which is why we extrapolate for 4000 qubits from the functional dependence in the range of $10^{3}..10^{18}$ T-gates. We then divide by $k=1.6$ to estimate the runtime of the complication that does not get rid of Cliffords but takes advantage of the parallelism. The results are:
Hubbard Ca3Co2O6
runtime (s) 38,000 $2.2\cdot 10^{17}$
($11$ hours) (7b years)
physical qubits 2.9M 46M
factory qubits 0.8M 4M
In the MIT material search, we need to execute $10^{3}$ runs for each model of the $10^{3}$ models. Even if we reduce their complexity to the same value as the Hubbard model above, one may still need to execute them in parallel on multiple quantum computers or take years as experts also need time to come up with each next model. We expect only the hardest models, say $1\%$ of 1000 suggested ones, to be sent to the quantum computer, while others would be investigated by classical methods or experimentally.